OpenStudy (anonymous):

What is the exact value of the square root of the quantity 63 over 16.? A:9 the square root of 7 all over 2. B:3 the square root of 7 all over 4. C:3 the square root of 21 all over 4. D:9 the square root of 7 all over 8.

ganeshie8 (ganeshie8):

factor 63 & 9 63 = 9 * 7 16 = 4 * 4

ganeshie8 (ganeshie8):